# Understanding Fine-tuning in Approximate Unlearning: A Theoretical Perspective

### v2

@@ B.1 Useful properties
See [30] for the proofs and for more properties. ∎
Corollary 1 (Projection Matrix properties).
Let $\mathbf{P}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Under 1, the remaining(forgetting) dataset matrix can be denoted as $\mathbf{X}_{r}=[\begin{array}[]{l}\mathbf{R}\\ \mathbf{0}\end{array}]$ and $\mathbf{X}_{f}=[\begin{array}[]{l}\mathbf{0}\\ \mathbf{F}\end{array}]$, where $\mathbf{R}\subseteq\mathbb{R}^{d_{r}\times{(n-n_{f})}}$ and $\mathbf{F}\subseteq\mathbb{R}^{d_{f}\times n_{f}}$ correspond to the non-zero parts. Then, it holds that:
@@
Corollary 2 (Minimum Norm Solution 1).
Let $\mathbf{P},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Then, the solution to the optimization problem Equation 1, Equation 2 and Equation 3 can be represented by:
@@
    Under 1, $\mathbf{w}_{o}=\mathbf{P}\mathbf{w}_{*}$, $\mathbf{w}_{t}=(\mathbf{I}-\mathbf{P}_{t})\mathbf{w}_{o}+\mathbf{P}_{t}\mathbf{w}_{*}^{r}$, and $\mathbf{w}_{g}=\mathbf{P}_{r}\mathbf{w}_{*}^{r}$;
@@
    Under 2, $\mathbf{w}_{o}=\mathbf{P}\mathbf{w}_{*}$, $\mathbf{w}_{t}=(\mathbf{I}-\mathbf{P}_{t})\mathbf{w}_{o}+\mathbf{P}_{t}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$, and $\mathbf{w}_{g}=\mathbf{P}_{r}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$;
@@ B.2 Proof of Theorem 1
[…]
where the second equality arises from the model setting and Proposition 2, while the penultimate equality is due to the properties of the projection matrix. According to Corollary 1, we have
[…]
@@
[…]
@@
[…]
@@ B.3 Proof of Theorem 2
Corollary 3 (Projection Matrix properties).
Let $\mathbf{P}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Under 2, it holds that:
@@
Proof of Corollary 3.
Proof of Corollary 3 follows the proof of Corollary 1 directly. ∎
Now we are ready to go through the proof of Theorem 2. Similar to the non-overlapping case, the golden model holds that
[…]
@@
where the second equality also arises from the model setting and Proposition 2, while the penultimate equality is due to the properties of the projection matrix. According to Corollary 3, we have
[…]
@@
[…]
@@
The last equality derives from that the facts the projection matrix is commutative matrix and the last property holds in Corollary 3. For the unlearning accuracy, it holds that
[…]
@@
where the penultimate equality is due to Corollary 3 and the last equality comes from the fact $\mathbf{X}_{t}$ enjoys the same data structure as $\mathbf{X}_{t}$ such that:
@@ B.4 Proof of Theorem 3
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ C.1 Useful properties
See (zarantonello1971projections) for the proofs and for more properties. ∎
Corollary C.1 (Projection Matrix properties).
Let $\mathbf{P}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Under ˜3.1, the remaining(forgetting) dataset matrix can be denoted as $\mathbf{X}_{r}=[\begin{array}[]{l}\mathbf{R}\\ \mathbf{0}\end{array}]$ and $\mathbf{X}_{f}=[\begin{array}[]{l}\mathbf{0}\\ \mathbf{F}\end{array}]$, where $\mathbf{R}\subseteq\mathbb{R}^{d_{r}\times{(n-n_{f})}}$ and $\mathbf{F}\subseteq\mathbb{R}^{d_{f}\times n_{f}}$ correspond to the non-zero parts. Then, it holds that:
@@
Corollary C.2 (Minimum Norm Solution 1).
Let $\mathbf{P},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Then, the solution to the optimization problem Equation˜1, Equation˜2 and Equation˜3 can be represented by:
@@
    Under ˜3.1, $\mathbf{w}_{o}=\mathbf{P}\mathbf{w}_{*}$, $\mathbf{w}_{t}=(\mathbf{I}-\mathbf{P}_{t})\mathbf{w}_{o}+\mathbf{P}_{t}\mathbf{w}_{*}^{r}$, and $\mathbf{w}_{g}=\mathbf{P}_{r}\mathbf{w}_{*}^{r}$;
@@
    Under ˜3.3, $\mathbf{w}_{o}=\mathbf{P}\mathbf{w}_{*}$, $\mathbf{w}_{t}=(\mathbf{I}-\mathbf{P}_{t})\mathbf{w}_{o}+\mathbf{P}_{t}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$, and $\mathbf{w}_{g}=\mathbf{P}_{r}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})$;
@@
$\text{RL:}\quad L(\mathbf{w}_{g},D_{r})$ $=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}_{g}-\mathbf{y}_{r}\|^{2}=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{P}_{r}\mathbf{w}_{*}^{r}-\mathbf{X}_{r}^{\top}\mathbf{w}_{*}^{r}\|^{2}=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}(\mathbf{P}_{r}-\mathbf{I})\mathbf{w}_{*}^{r}\|^{2}=0,$
where the second equality arises from the model setting and Proposition C.2, while the penultimate equality is due to the properties of the projection matrix. According to Corollary˜C.1, we have
$\text{UL:}\quad L(\mathbf{w}_{g},D_{f})$ $=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}_{g}-\mathbf{y}_{f}\|^{2}=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{P}_{r}\mathbf{w}_{*}^{r}-\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}\|^{2}$
@@
$\text{RL:}\quad L(\mathbf{w}_{t},D_{r})$ $=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}_{t}-\mathbf{y}_{r}\|^{2}=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}((\mathbf{I}-\mathbf{P}_{t})\mathbf{w}_{o}+\mathbf{P}_{t}\mathbf{w}_{*}^{r})-\mathbf{X}_{r}^{\top}\mathbf{w}_{*}^{r}\|^{2}$
@@
$\text{UL:}\quad L(\mathbf{w}_{t},D_{f})$ $=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}_{t}-\mathbf{y}_{f}\|^{2}=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}((\mathbf{I}-\mathbf{P}_{t})\mathbf{w}_{o}+\mathbf{P}_{t}\mathbf{w}_{*}^{r})-\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}\|^{2}$
@@
Corollary C.3 (Projection Matrix properties).
Let $\mathbf{P}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top},\mathbf{P}_{r},\mathbf{P}_{f},\mathbf{P}_{t}$ be the corresponding projection operator for $\mathbf{X},\mathbf{X}_{r},\mathbf{X}_{f},\mathbf{X}_{t}$ respectively. Under ˜3.3, it holds that:
@@
Proof of Corollary˜C.3.
Proof of Corollary˜C.3 follows the proof of Corollary˜C.1 directly. ∎
Now we are ready to go through the proof of Theorem˜3.4. Similar to the non-overlapping case, the golden model holds that
$\text{RL:}\quad L(\mathbf{w}_{g},D_{r})$ $=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}_{g}-\mathbf{y}_{r}\|^{2}=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{P}_{r}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})-\mathbf{X}_{r}^{\top}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})\|^{2}$
@@
where the second equality also arises from the model setting and Proposition C.2, while the penultimate equality is due to the properties of the projection matrix. According to Corollary˜C.3, we have
$\text{UL:}\quad L(\mathbf{w}_{g},D_{f})$ $=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}_{g}-\mathbf{y}_{f}\|^{2}=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{P}_{r}(\mathbf{w}_{*}^{r}+\mathbf{w}_{*}^{lap})-\mathbf{X}_{f}^{\top}(\mathbf{w}_{*}^{f}+\mathbf{w}_{*}^{lap})\|^{2}$
@@
$\text{RL:}\quad L(\mathbf{w}_{t},D_{r})$ $=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}_{t}-\mathbf{y}_{r}\|^{2}$
@@
The last equality derives from that the facts the projection matrix is commutative matrix and the last property holds in Corollary˜C.3. For the unlearning accuracy, it holds that
$\text{UL:}\quad L($ $\mathbf{w}_{t},D_{f})=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}_{t}-\mathbf{y}_{f}\|^{2}$
@@
where the penultimate equality is due to Corollary˜C.3 and the last equality comes from the fact $\mathbf{X}_{t}$ enjoys the same data structure as $\mathbf{X}_{t}$ such that:
@@
$\text{RL:}\quad L(\mathbf{w}_{t},D_{r})$ $=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}_{t}-\mathbf{y}_{r}\|^{2}=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}((\mathbf{I}-\mathbf{P}_{t})\hat{\mathbf{w}}_{o}+\mathbf{P}_{t}\mathbf{w}_{*}^{r})-\mathbf{X}_{r}^{\top}\mathbf{w}_{*}^{r}\|^{2}$
@@
$\text{UL:}\quad L(\mathbf{w}_{t},D_{f})$ $=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}_{t}-\mathbf{y}_{f}\|^{2}=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}((\mathbf{I}-\mathbf{P}_{t})\hat{\mathbf{w}_{o}}+\mathbf{P}_{t}\mathbf{w}_{*}^{r})-\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}\|^{2}$
@@
$\text{RL:}\quad L(\mathbf{w}_{t},D_{r})$ $=\frac{1}{n_{r}}\|\mathbf{X}_{r}^{\top}\mathbf{w}_{t}-\mathbf{y}_{r}\|^{2}$
@@
$\text{UL:}\quad L(\mathbf{w}_{t},D_{f})$ $=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{w}_{t}-\mathbf{y}_{f}\|^{2}$